1. The area region bounded by the parabolas y2=4ax and x2=4ay is

Grade:12th pass

1 Answers

Arun
6 years ago
We have, y2 = 4ax --------------------------- (1) 
x2 = 4ay ---------------------------- (2) 
(1) and (2) intersects hence 
x = y2/4a (a > 0) 
=> (y2/4a)2 = 4ay 
=> y4 = 64a3
=> y4 – 64a3y = 0 
=> y[y3 – (4a)3] = 0 
=> y = 0, 4a 
When y = 0, x = 0 and when y = 4a, x = 4a. 
The points of intersection of (1) and (2) are O(0, 0) and A(4a, 4a). 
The area of the region between the two curves 
= Area of the shaded region 
04a(y1 – y2)dx 
04a[√(4ax) – x2/4a]dx 
= [2√a.(x3/2)/(3/2) – (1/4a)(x3/3)]04a 
= 4/3√a(4a)3/2 – (1/12a)(4a)3 – 0 
= 32/3a2 – 16/3a2 
= 16/3a2 sq. units